# Quantum-enhanced Markov Chain Monte Carlo for systems larger than your Quantum Computer

### v1

@@ Abstract
Quantum computers theoretically promise computational advantage in many tasks, but it is much less clear how such advantage can be maintained when using existing and near-term hardware that has limitations in the number and quality of its qubits. One promising application was proposed in Layden et al [1] where a method to reduce the thermalisation time required when sampling from hard probability distribution was introduced as a Quantum-enhanced Markov Chain Monte Carlo (QeMCMC) approach. In [1] the size of the required quantum computer scales linearly with the problem, putting limitations on the sizes of systems that one can consider. In this work we introduce a framework to coarse grain the algorithm in such a way that the quantum computation can be performed using, multiple times, smaller quantum computers and we term the method the Coarse Grained Quantum-enhanced Markov Chain Monte Carlo (CGQeMCMC). Example strategies within this framework are put to the test, with the quantum speedup of [1] persisting while using only $\sqrt{n}$ simulated qubits where $n$ is the number of qubits required in the original QeMCMC – a quadratic reduction in resources. The coarse graining framework has the potential to be practically applicable in the near term as it requires very few qubits to approach classically intractable problem instances, here only 6 simulated qubits suffice to gain advantage compared to standard classical approaches when investigating the magnetisation of a 36 spin system. Our method is also adjustable to quantum hardware specifications, and it appears that it can be easily combined with other techniques both classical and quantum.
[…]
[first]organization=Quantum Software Lab, School of Informatics, addressline=The University of Edinburgh, city=Edinburgh, country=United Kingdom
1 Introduction
@@
Sampling from hard distributions is a problem ubiquitous in science, technology, finance and statistics. While there are well understood families of traditional computational methods such as the Markov Chain Monte Carlo (MCMC) [2, 3] that attempt to address this challenge, the advent of quantum computing offers a promising avenue for considerably more efficient solutions in randomised algorithms such as [4] and [5].
@@
Current Quantum Processing Units (QPUs) are labelled as ‘Noisy Intermediate-Scale Quantum’ (NISQ) devices, meaning that although they exhibit remarkable quantum phenomena they have strict limitations in the number of available qubits and the quality of qubits and operations that are very error-prone. In this era where full quantum error-correction is out of reach, one must explore algorithms that are not only robust to noise but also as efficient in the number of qubits as possible. One algorithm that addresses the former requirement is the Quantum-enhanced Markov Chain Monte Carlo (QeMCMC) of [1]. It introduces an elegant way to enhance the classical MCMC by “quantising” one part of the classical MCMC algorithm: namely the proposal step. The method has also been extended, by application to Hamiltonian Monte-Carlo [6] and Variational algorithms [7].
@@
In this paper, we propose a way to significantly reduce the qubit requirement of the QeMCMC, while maintaining quantum speedup over the classical MCMC in the low temperature limit [^3] Here we need to stress that by quantum speed up we mean that our approach outperforms the classical MCMC. It is understood, that while the size of the quantum computer used is small, and specifically smaller than what can be classically simulated, the advantage of our algorithm can be viewed as a “quantum inspired” approach, while the true quantum advantage can be achieved when the QPU used is larger than what can be classically simluated.. This is done by Coarse Graining (CG) the problem instance - an Ising model lattice - in such a way that each subset “group” of the lattice can be evaluated in separate quantum computations, even on different QPU’s in parallel. In this way, it is predicted that one can use a QPU with as few as $O(\sqrt{n})$ qubits and still achieve quantum advantage. We call this method the Coarse Grained Quantum–enhanced Markov Chain Monte Carlo (CGQeMCMC).
@@
The way in which QeMCMC [1] achieves quantum advantage is by reducing the thermalisation time of a Markov chain. This is the number of steps taken to reach the stationary distribution which you wish to sample from. Here we need to make a distinction between the cost of proving that thermalisation is achieved fast (a task needed to prove the performance of the algorithm) and the use of our CGQeMCMC practically for some task. For the former, the primary method of analysing thermalisation is by spectral gap analysis, a well known and unambiguous figure of merit. Calculating the spectral gap, however, is an entirely different process from MCMC and is exponentially more expensive, restricting our spectral gap analysis to $10$ spins. For the latter, we use Markov chains to estimate an example observable - magnetisation - and in this case we consider a system of $36$ spins to experimentally witness quick thermalisation of the CGQeMCMC. In both cases, the quantum computers part of CGQeMCMC is realised in an emulation environment (using classical devices – see A for details of the implementation).
@@
  • 1.
    Propose a framework for applying course graining to the QeMCMC with the Ising model used as the test case. The framework allows the problem instance to be altered to reduce the number of qubits required and, in some cases, parallelise computation over multiple QPU.
  • 2.
@@
  • 3.
    Simulation of CGQeMCMC for larger Ising model instances up to 36 spins through the calculation of average magnetisation and analysis of convergence.
  • 4.
@@
The paper is organised as follows: In Section 2 we give the necessary background namely, the classical MCMC algorithm, the spin glass Ising model, the problem of thermalisation and the QeMCMC of [1]. In Section 3 we propose a general framework for coarse graining the Ising model so that smaller quantum computers can be used to tackle larger instances within the CGQeMCM and even achieve quantum advantage. We also discuss figures of merit of simple Markov chains and the difficulties in assessing thermalisation. Section 4 discusses the particular coarse grained proposal methods employed, before they are explored numerically in Section 5. Conclusion and future directions are given in Section 6.
2 Preliminaries
@@
In this section, we introduce the classical Ising model, provide an overview of MCMC and the issue of thermalisation for difficult problem instances. We then introduce the QeMCMC which improves the thermalisation by utilizing quantum computation.
2.1 The Ising model
@@
The archetypal test-case of random sampling algorithms such as MCMC are “spin glass” Ising models. “Glassy” due to their magnetic disorder in analogy with the positional disorder in chemical glass, such lattices are composed of spins that can exist in either states $s_{i}=\{1,-1\}$. The state, $\boldsymbol{s}$, composed of n spins is thus defined $\boldsymbol{s}=(s_{1},s_{2}...s_{n})$. Each configuration of spins $\{1,-1\}^{n}$ has an associated energy ($E(\boldsymbol{s})$), Boltzmann probability ($\mu(\boldsymbol{s})$) and magnetisation ($m(\boldsymbol{s})$) defined below:
$E(\boldsymbol{s})=-\sum_{j>k=1}^{n}J_{jk}s_{j}s_{k}-\sum_{j=1}^{n}h_{j}s_{j}$ (1)
$\mu(\boldsymbol{s})=\frac{1}{Z}e^{-E(\boldsymbol{s})/T}$ (2)
$m(\boldsymbol{s})=\frac{1}{n}\sum_{j=1}^{n}s_{j}$ (3)
@@
where $h$ and $J$ are the field and coupling respectively that define the Ising model instance, $Z$ is the partition function $Z=\sum_{\boldsymbol{s}}e^{-E(\boldsymbol{s})/T}$ and $T$ is a temperature parameter. The Boltzmann average magnetisation $\langle m\rangle_{\mu}=\sum_{\boldsymbol{s}}\mu(\boldsymbol{s})m(\boldsymbol{s})$ is an example of an observable which can be hard to calculate due to the size of the state space. For just 100 spins, the number of possible configurations are $2^{100}\approx 10^{30}$ and requiring an algorithm to sum over these becomes intractable.
@@
For systems where $J$ and $h$ are unstructured and have values of varying signs, the energy landscape is particularly rugged and hard to explore. When one wishes compute the value of an observable from this landscape, in the low temperature limit, very few states of low energy contribute. These states are often very far apart in Hamming distance  [^4] Here, Hamming distance between two states is defined as the number of positions that the two states have different spin values., meaning that transitioning between them proportionally to their Boltzmann probability is extremely inefficient.
2.2 Markov Chain Monte Carlo
@@
The MCMC has a remarkable history. Named as one of the ten most influential algorithms in history, it was developed to calculate observables of a probability distribution that is too highly dimensional to be evaluated analytically or with brute force numerics [8]. The first component is the Markov chain - a walk on such a highly dimensional landscape - where the probability of each step only depends on the last position. The chain is normally produced by the Metropolis-Hastings algorithm which, with some probability (hence “Monte-Carlo”), rejects or accepts update steps based on the value of the probability distribution at that point [9, 10].
@@
MCMC is able to skirt the responsibility of ever directly calculating the expensive (due to calculation of $Z$) $\mu(s)$. Instead, it performs an approximation by exploring the state space by random walk from an arbitrary starting point. Each step in the algorithm is composed of two sub-steps: proposal and acceptance. The proposal step determines the probability of state $\boldsymbol{s}^{\prime}$ being proposed when the system is in state $\boldsymbol{s}$. It is represented by the matrix $Q(\boldsymbol{s}^{\prime}|\boldsymbol{s})$ and can be constructed in many different ways including through the use of a quantum computer as we see in Section 2.4. The acceptance step is often governed by Metropolis-Hastings (MH), where the matrix $A(\boldsymbol{s}^{\prime}|\boldsymbol{s})$ (Equation 4) dictates the probability that a move is accepted [11, 9].
$A\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)=\min\left(1,\frac{\mu\left(\boldsymbol{s}^{\prime}\right)}{\mu(\boldsymbol{s})}\frac{Q\left(\boldsymbol{s}\mid\boldsymbol{s}^{\prime}\right)}{Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)}\right)$ (4)
@@
// Initialisation
2 $\boldsymbol{s}\leftarrow$ initial spin configuration;
@@ 5      $s^{\prime}\leftarrow$ update proposal;
7      $A\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)=\min\left(1,\frac{\mu\left(\boldsymbol{s}^{\prime}\right)}{\mu(\boldsymbol{s})}\frac{Q\left(\boldsymbol{s}\mid\boldsymbol{s}^{\prime}\right)}{Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)}\right)$
@@
9           $\boldsymbol{s}=\boldsymbol{s}^{\prime}$;
@@
The importance of the acceptance step underlies any future algorithm. One does not need to calculate $\mu(\boldsymbol{s})$ or $\mu(\boldsymbol{s}^{\prime})$ at each step, merely their ratio which can be computed in $O\left(n^{2}\right)$ polynomial time [9]. Together, proposal and acceptance steps result in the probability matrix, $P\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)=A\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)Q\left(\boldsymbol{s}^{\prime}\mid\boldsymbol{s}\right)$, which is the probability that each step in the algorithm will result in transition from $\boldsymbol{s}$ to $\boldsymbol{s}^{\prime}$.
@@
To ensure convergence, the Markov chain must be irreducible and aperiodic meaning that any state can be (eventually) reached from any other state and that there exists no repeating loops in the chain. Detailed balance ensures both by requiring the following is satisfied for any transition $\boldsymbol{s}\rightarrow\boldsymbol{s}^{\prime}$.
$\mu(\boldsymbol{s})P(\boldsymbol{s}|\boldsymbol{s}^{\prime})=\mu(\boldsymbol{s}^{\prime})P(\boldsymbol{s}^{\prime}|\boldsymbol{s})$ (5)
@@
In this context, the probability distribution $\mu(\boldsymbol{s})$ is often labeled as the “stationary” or “target” distribution. When the temperature drops the exponential term in the Boltzmann distribution becomes increasingly sensitive to small differences in energy levels. Consequently, the acceptance probability for proposed transitions becomes negligible, leading to extremely slow exploration of the state space. This phenomenon impedes efficient sampling and hinders convergence, as only proposals of similar or lower energy are consistently accepted. This means the Markov-chain in a relatively low energy state lies dormant until it happens across one of only a few low energy states, which takes exponential time. The QeMCMC aims to address this issue as explored in Section 2.4.
2.3 Thermalisation

### v2

@@ Abstract
Quantum computers theoretically promise computational advantage in many tasks, but it is much less clear how such advantage can be maintained when using existing and near-term hardware that has limitations in the number and quality of its qubits. Layden et al. [Nature 619, 282 (2023)] proposed a promising application by introducing a Quantum-enhanced Markov Chain Monte Carlo (QeMCMC) approach to reduce the thermalization time required when sampling from hard probability distributions. In QeMCMC the size of the required quantum computer scales linearly with the problem, putting limitations on the sizes of systems that one can consider. In this work we introduce a framework to coarse grain the algorithm in such a way that the quantum computation can be performed using considerably smaller quantum computers and we term the method the Coarse Grained Quantum-enhanced Markov Chain Monte Carlo (CGQeMCMC). Example strategies within this framework are put to the test, with the quantum speedup persisting while using only $\sqrt{n}$ simulated qubits where $n$ is the number of qubits required in the original QeMCMC – a quadratic reduction in resources. The coarse graining framework has the potential to be practically applicable in the near term as it requires very few qubits to approach classically intractable problem instances; in this case only 6 simulated qubits suffice to gain advantage compared to standard classical approaches when investigating the magnetization of a 36 spin system. Our method can be easily combined with other classical and quantum techniques and is adaptable to various quantum hardware specifications - in particular those with limited connectivity.
preprint: APS/123-QED
I Introduction
@@
Sampling from hard distributions is a problem ubiquitous in science, technology, finance and statistics. While there are well understood families of traditional computational methods such as the Markov Chain Monte Carlo (MCMC) Geyer (1992); Duane et al. (1987); Brooks et al. (2011) that attempt to address this challenge, the advent of quantum computing offers a promising avenue for considerably more efficient solutions in randomized algorithms. This includes algorithms for Monte Carlo Integration (MCI) Herbert (2022); Akhalwaya et al. (2023), Quantum Metropolis Sampling (QMS) Temme et al. (2011); Yung and Aspuru-Guzik (2012) as well as a very general quantum speedup of Monte Carlo methods Montanaro (2015).
@@
Current Quantum Processing Units (QPUs) are labeled as ‘Noisy Intermediate-Scale Quantum’ (NISQ) devices, meaning that although they exhibit remarkable quantum phenomena they have strict limitations in the number and quality of available qubits. In this era where full quantum error-correction is out of reach, one must explore algorithms that are not only robust to noise but are also efficient in the number of qubits required. One algorithm that addresses the former requirement is the Quantum-enhanced Markov Chain Monte Carlo (QeMCMC) of Layden et al. (2023). It introduces an elegant way to enhance the classical MCMC by “quantizing” one part of the classical MCMC algorithm: namely the proposal step.
The QeMCMC has some promising properties alongside its noise resilience. It is expected to exceed quadratic scaling in comparison with the best classical alternative in certain contexts, making it particularly promising for early fault-tolerant quantum processors. Additionally, QeMCMC functions as a one-shot algorithm, requiring only a single execution on a QPU per step in the hybrid algorithm by avoiding the need for expectation value estimation or state tomography. The method has also been extended, by application to Hamiltonian Monte Carlo Lockwood et al. (2024) and Variational algorithms Nakano et al. (2023).
@@
In this paper, we propose a way to significantly reduce the qubit requirement of the QeMCMC, while maintaining a scaling advantage that may lead to quantum speedup over the classical MCMC in the low temperature limit  [^1] Here we need to stress that by quantum speed up we mean that our approach outperforms the classical MCMC. It is understood, that while the size of the quantum computer used is small, and specifically smaller than what can be classically simulated, the advantage of our algorithm can be viewed as a “quantum inspired” approach, while the true quantum advantage can be achieved when the QPU used is larger than what can be classically simluated.. This is done by Coarse Graining (CG) the problem instance - an Ising model lattice - in such a way that each subset “group” of the lattice can be evaluated in separate quantum computations, even on different QPUs in parallel. In this way, it is predicted that one can use a QPU with as few as $O(\sqrt{n})$ (where n is the number of spins in the Ising model) qubits and still achieve a scaling advantage in comparison with classical techniques. We call this method the Coarse Grained Quantum–enhanced Markov Chain Monte Carlo (CGQeMCMC).
@@
The way in which the QeMCMC achieves a scaling advantage is by reducing the thermalization time of a Markov chain. This is the number of steps taken to reach the stationary distribution which you wish to sample from. Here we need to make a distinction between the cost of proving that thermalization is achieved fast (a task needed to prove the performance of the algorithm) and the use of our CGQeMCMC practically for some task. For the former, the primary method of analysing thermalization is by spectral gap analysis, a well known and unambiguous figure of merit. Calculating the spectral gap, however, is an entirely different process from MCMC and is exponentially more expensive, restricting our spectral gap analysis to $10$ spins. For the latter, we use Markov chains to estimate an example observable - magnetization - and in this case we consider a system of $36$ spins to experimentally witness quick thermalization of the CGQeMCMC. In both cases, the quantum computers part of CGQeMCMC is realised in an emulation environment (using classical devices – see Appendix A for details of the implementation).
@@
  • Propose a framework for applying course graining to the QeMCMC with the Ising model used as the test case. The framework allows the problem instance to be altered to reduce the number of qubits required and, in some cases, parallelize computation over multiple QPU.
@@
  • Simulation of CGQeMCMC for larger Ising model instances up to 36 spins through the calculation of average magnetization and analysis of convergence.
@@
The paper is organized as follows: In Section II, we give the necessary background, namely, the classical MCMC algorithm, the spin glass Ising model, the problem of thermalization and the QeMCMC of Layden et al. (2023). In Section III we propose the CGQeMCMC, a general framework for coarse graining the Ising model so that smaller quantum computers can be used to tackle larger problem instances. We also discuss the figures of merit of simple Markov chains and the difficulties in assessing thermalization. Section IV discusses the particular coarse grained proposal methods employed, before they are explored numerically in Section V. Conclusion and future directions are given in Section VI.
II Preliminaries
@@
In this section, we introduce the classical Ising model, provide an overview of MCMC and the issue of thermalization for difficult problem instances. We then introduce the QeMCMC which improves the thermalization by utilizing quantum computation.
II.1 The Ising model
@@
The archetypal test-case of random sampling algorithms such as MCMC is the “spin glass” Ising model. Analyzing these systems is crucial due to their wide-ranging applications in fields such as materials science, biology (protein folding), and computer science (artificial neural networks). “Glassy” due to their magnetic disorder in analogy with the positional disorder in chemical glass, such lattices are composed of spins that can exist in either states $s_{i}=\{+1,-1\}$. The state $\mathbf{s}$, composed of n spins, is thus defined $\mathbf{s}=(s_{1},s_{2}...s_{n})$. Each configuration of spins $\{1,-1\}^{n}$ has an associated energy ($E(\mathbf{s})$), Boltzmann probability ($\mu(\mathbf{s})$), and magnetization ($m(\mathbf{s})$) defined below:
$E(\mathbf{s})=-\sum_{j>k=1}^{n}J_{jk}s_{j}s_{k}-\sum_{j=1}^{n}h_{j}s_{j}$ (1)
$\mu(\mathbf{s})=\frac{1}{Z}e^{-E(\mathbf{s})/T}$ (2)
$m(\mathbf{s})=\frac{1}{n}\sum_{j=1}^{n}s_{j}$ (3)
@@
Where $h$ and $J$ are the field and coupling respectively that define the Ising model instance, $Z$ is the partition function $Z=\sum_{\mathbf{s}}e^{-E(\mathbf{s})/T}$ and $T$ is a temperature parameter. The Boltzmann average magnetization $\langle m\rangle_{\mu}=\sum_{\mathbf{s}}\mu(\mathbf{s})m(\mathbf{s})$ is an example of an observable which can be hard to calculate due to the size of the state space. For just 100 spins, the number of possible configurations are $2^{100}\approx 10^{30}$ and requiring an algorithm to sum over these becomes intractable.
@@
For systems where $J$ and $h$ are unstructured and have values of varying signs, the energy landscape is particularly rugged and hard to explore. When one wishes compute the value of an observable from this landscape, in the low temperature limit, very few states of low energy contribute. Sampling from these contributing states is particularly difficult as they can be obscured from each other by large hamming distance  [^2] Here, Hamming distance between two states is defined as the number of positions that the two states have different spin values. and cost barriers. Thus, when in one state transitioning to another is typically very hard.
II.2 Markov Chain Monte Carlo
@@
The MCMC has a remarkable history. Named as one of the ten most influential algorithms in history, it was developed to calculate observables of a probability distribution that is too highly dimensional to be evaluated analytically or with brute force numerics Dongarra and Sullivan (2000). The first component is the Markov chain - a walk on such a highly dimensional landscape - where the probability of each step only depends on the last position. The chain is normally produced by the Metropolis-Hastings algorithm which, with some probability (hence “Monte Carlo”), rejects or accepts update steps based on the value of the probability distribution at that point Metropolis et al. (1953); Fishman (2013).
@@
MCMC is able to skirt the responsibility of ever directly calculating the expensive (due to calculation of $Z$) $\mu(s)$. Instead, it performs an approximation by exploring the state space by random walk from an arbitrary starting point. Each step in the algorithm is composed of two sub-steps: proposal and acceptance. The proposal step determines the probability of state $\mathbf{s}^{\prime}$ being proposed when the system is in state $\mathbf{s}$. It is represented by the matrix $Q(\mathbf{s}^{\prime}|\mathbf{s})$ and can be constructed in many different ways including through the use of a quantum computer as we see in Section II.4. The acceptance step is often governed by Metropolis-Hastings (MH), where the matrix $A(\mathbf{s}^{\prime}|\mathbf{s})$ (Equation 4) dictates the probability that a move is accepted Hastings (1970); Metropolis et al. (1953).
$A\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)=\min\left(1,\frac{\mu\left(\mathbf{s}^{\prime}\right)}{\mu(\mathbf{s})}\frac{Q\left(\mathbf{s}\mid\mathbf{s}^{\prime}\right)}{Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)}\right)$ (4)
@@
// Initialization
2 $\mathbf{s}\leftarrow$ initial spin configuration;
@@ 5      $s^{\prime}\leftarrow$ update proposal;
7      $A\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)\leftarrow\min\left(1,\frac{\mu\left(\mathbf{s}^{\prime}\right)}{\mu(\mathbf{s})}\frac{Q\left(\mathbf{s}\mid\mathbf{s}^{\prime}\right)}{Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)}\right)$
@@
9           $\mathbf{s}\leftarrow\mathbf{s}^{\prime}$;
@@
The importance of the acceptance step underlies any future algorithm. One does not need to calculate $\mu(\mathbf{s})$ or $\mu(\mathbf{s}^{\prime})$ at each step, merely their ratio which can be computed in $O\left(n^{2}\right)$ polynomial time Metropolis et al. (1953). Together, proposal and acceptance steps result in the probability matrix, $P\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)=A\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)Q\left(\mathbf{s}^{\prime}\mid\mathbf{s}\right)$, which is the probability that each step in the algorithm will result in transition from $\mathbf{s}$ to $\mathbf{s}^{\prime}$.
@@
To ensure convergence, the Markov chain must be irreducible and aperiodic meaning that any state can be (eventually) reached from any other state and that there exists no repeating loops in the chain. These are true in our case as no states are forbidden assuming there is no unwanted symmetry as discussed in Layden et al. (2023). Detailed balance is another important condition that ensures the stationary distribution of our Markov Chain is the probability distribution from which we wish to sample. It requires that the following is satisfied for any transition $\mathbf{s}\rightarrow\mathbf{s}^{\prime}$ where $s\neq s^{\prime}$.
$P(\mathbf{s}^{\prime}|\mathbf{s})\mu(\mathbf{s})=P(\mathbf{s}|\mathbf{s}^{\prime})\mu(\mathbf{s^{\prime}})$ (5)
@@
In this context, the probability distribution $\mu(\mathbf{s})$ is often labeled as the “stationary” or “target” distribution. When the temperature drops the exponential term in the Boltzmann distribution becomes increasingly sensitive to small differences in energy levels. Consequently, the acceptance probability for proposed transitions becomes negligible, leading to extremely slow exploration of the state space. This phenomenon impedes efficient sampling and hinders convergence, as only proposals of similar or lower energy are consistently accepted. This means the Markov-chain in a relatively low energy state lies dormant until it happens across one of only a few low energy states, which takes exponential time. The QeMCMC aims to address this issue as explored in Section II.4.
II.3 Thermalization